OpenStudy (anonymous):

what is the exact value of cos60sin60tan60

OpenStudy (p0sitr0n):

cos60 = cos(pi/3) = 1/2 sin60 = sin(pi/3) = sqrt3/2 tan60 = sin/cos = sqrt3/2 * 2 = sqrt 3 1/2 * sqrt3/2 * sqrt3 = 1/4 * 3 = 3/4

OpenStudy (rajee_sam):

Sin^2 (60)